OpenStudy (anonymous):

Find the slope of the line connecting these points. (2, 9), (-4, 4)

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

i dont get it :/

hartnn (hartnn):

ok,first, whats the difference of y-co-ordinates ? 9-4 =.. ?

hartnn (hartnn):

you there ?

hartnn (hartnn):

then you need to take difference in x-co-ordinates 2- (-4) =... ? then just divide them and you get the slope! :)

OpenStudy (anonymous):

ok thanks